# Five-loop Anomalous Dimensions of Cubic Scalar Theory from Operator Product Expansion

### v2

@@ 3 Two-loop demonstration
For illustrative purposes, in this section we will present detailed computations of the two-loop anomalous dimensions of the $\varphi^{Q}$ operator. This requires computing the field renormalization $Z_{\varphi}$ and $Z_{\sigma}$-factors, the coupling renormalization $Z_{h}$ and $Z_{g}$-factors and the $\varphi^{Q}$ operator renormalization $Z_{\varphi^{Q}}$-factor. These factors are determined from the corresponding two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions respectively. All results related to Feynman diagrams are computed by private Mathematica code, while the UV divergences of two-point integrals are computed by HyperlogProcedures Maple package HP . The strategy of computation is formalized as a sequential algorithm, following the computations of the symmetry factors of topologies, the coupling factors of Feynman diagrams, the UV divergence of two-point integrals and finally the determination of $Z$-factors from UV finiteness conditions.
@@ 3.1 Two-point correlation functions and the field renormalization
Let us begin with the field renormalization computation. The $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ can be computed from the two-point correlation functions $\langle\varphi\bar{\varphi}\rangle$ and $\langle\sigma\sigma\rangle$ respectively as in eqn.(21). The one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(2).
@@
Each topology represents a collection of diagrams with distinct allowable internal field configurations of $\sigma,\varphi_{i}$, and these collections are characterized by coupling factors. The symmetry factors and the coupling factors of these topologies are listed in the table below,
@@
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$ 1/2 $h_{0}^{2}+Ng_{0}^{2}$ $2g_{0}^{2}$
$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$ 1/2 $N(2g_{0}^{4}+g_{0}^{2}h_{0}^{2})+h_{0}^{4}$ $Ng_{0}^{4}+2g_{0}^{4}+g_{0}^{2}h_{0}^{2}$
$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ 1/2 $N(g_{0}^{4}+2g_{0}^{3}h_{0})+h_{0}^{4}$ $2g_{0}^{4}+2g_{0}^{3}h_{0}$
There is no need to cut any external legs of these topologies, and their corresponding UV divergences can be directly computed by TwoPoint command in HyperlogProcedures. Feynman integrals in the $\overline{\mbox{MS}}$ dimensional regularization scheme can be evaluated as power series of parameter $\epsilon$ with the results(an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown),
@@
Then the bare correlation functions are readily obtained as,
@@
where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factor. Attention should be paid to the $\pm$ signs of individual terms in the summation. The bare correlation function is computed by summing contributions from all tree-level and loop Feynman diagrams. Note that the TwoPoint command only computes the scalar integrals, omitting all $i$ factors from, for instance, the vertices or propagators. When summing these contributions, we must recover the factor $i$ for each loop order. Each vertex carries a factor of i-i. Each propagator contains a factor of $i$ in its numerator, but after Wick rotation to Euclidean space, the denominator acquires an additional factor of 1-1, resulting in an overall factor of i-i for each propagator. Each loop integration gains a factor of $i$ from the Wick rotation $\ell_{0}\to i\ell_{E0}$. Supposing a loop graph has $n_{v}$ vertices, $n_{e}$ propagators, $n_{\ell}$ loops and $n_{p}$ external legs, then the total factor would be $(-i)^{n_{v}}(-i)^{n_{e}}i^{n_{\ell}}$. From graph theory, we have the relations $n_{v}+n_{\ell}-n_{e}=1$ and $3n_{v}=2n_{e}+n_{p}$, and after calculation we get the final factor $(-1)^{n_{p}}i$ for each loop order. For two-point functions, $n_{p}=2$, so each loop order has a factor $i$, whereas the tree-level contrition has a factor i-i. If we factor out an overall i-i, the two-point correlation function can be evaluated as the tree-level contribution minus the sum of loop contributions.
@@
The renormalized two-point correlation functions eqn.(21) should be UV finite, so we get the UV finiteness conditions
$Z_{\varphi}G(g_{0},h_{0};\varphi(x_{0}),\bar{\varphi}(x_{1}))=\mbox{UV~finite}~~~,~~~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{0}),\sigma(x_{1}))=\mbox{UV~finite}~,$ (42)
@@
which means they should be free of $\epsilon$ poles. In order to get the equations of UV finiteness conditions, we rewrite the bare couplings $g_{0},h_{0}$ in the bare correlation functions to $g,h$ by relations eqn.(7), then expand the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$. Note that in the rewriting of couplings with eqn.(7), each $g$ or $h$ is accompanied by a factor $\mu^{\frac{\epsilon}{2}}$, so from the coupling factors we know that at $L$-loop order there will be a $\mu^{L\epsilon}$ factor. Combined with the $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor, at $L$-loop order there will be an overall $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor. Expansion of the factor around $\epsilon\to 0$ produces powers of $\ln\frac{\mu^{2}}{p^{2}}$ terms. Since each $\ln\frac{\mu^{2}}{p^{2}}$ order is independent, the UV finiteness conditions also require the coefficient of each $\ln\frac{\mu^{2}}{p^{2}}$ order is free of $\epsilon$ poles. Hence the higher Log order terms will produce additional equations. Inspecting the UV divergence structures of $L$-loop integrals, the leading divergence is $1/\epsilon^{L}$}, thus at the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ order there will in general be $L$ equations for the cancellation of $L$ poles. At the $\ln^{k}\frac{\mu^{2}}{p^{2}}$ order, the leading divergence is $1/\epsilon^{L-k}$, so there will be (Lk)(L-k) equations. We can either use the equations from the leading Log terms [^3] The leading Log term denotes terms with $\ln^{0}(\mu^{2}/p^{2})$ dependence, and the sub-Log term denotes terms with $\ln(\mu^{2}/p^{2})$ dependence, etc., or include those equations from higher Log terms. If we only use equations of the leading Log terms, then all equations of $Z_{\varphi},Z_{\sigma},Z_{g},Z_{h}$ should be solved simultaneously to completely determine the $Z$-factors. However, if we also use the equations of sub-Log terms, then each $Z$-factor can be determined by solving its own equations of UV finiteness conditions. This will be demonstrated in the following computations.
@@
Now let us take a look at the UV finiteness conditions. For $Z_{\varphi}$-factor, at one-loop, after the above prescribed expansions, we get
@@
The coefficients of sub-Log terms are already UV finite, and they will not provide any additional equations. In order for the leading Log term to be finite, the $1/\epsilon$ pole should cancel, and we get
@@
At two-loop order, we need to solve $z^{21}_{\varphi},z^{22}_{\varphi}$. Again after expansions, the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles at leading Log term produces two equations,
@@
and the vanishing of $1/\epsilon$ pole at sub-Log term produces one equation,
@@
In the traditional approach, the renormalization computation should be performed order by order. This is similar to the situation that solving the $Z$-factors using only equations of leading Log terms. For instance, by using two equations of leading Log term shown above, the determination of $z_{\varphi}^{21},z_{\varphi}^{22}$ will also rely on the results of all one-loop $Z$-factors, including $Z_{g},Z_{h},Z_{\varphi},Z_{\sigma}$, which must be determined a priori. Alternatively, if considering the UV finiteness conditions of all Log terms of a correlation function, the generated equations are sufficient to completely solve the corresponding $Z$-factor. This means that the computation does not rely on other $Z$-factors of lower-loop orders. For instance, the equation from sub-Log term shown above can be solved to obtain a combination of one-loop $Z$-factors. This one-loop $Z$-factor combination is exactly the one appearing in the equations of leading Log term, and after substituting it back to those two equations, we can completely determine the two-loop $Z_{\varphi}$-factor. Furthermore, as a by-product, the equations of sub-Log terms provide constraints to other $Z$-factors at lower-loop orders, which can be served as consistent checks for lower-loop $Z$-factors computed by other correlation functions.
Consequently, we determine the two-loop $Z_{\varphi}$-factors as,
@@
The determination of $Z_{\sigma}$-factor follows exactly the same procedures, and the only difference is the coupling factors. After the same computations, eventually we determine the two-loop $Z_{\sigma}$-factor as
@@
The higher-loop $Z_{\varphi},Z_{\sigma}$-factors can be determined following the same procedures. However there would be much more Feynman graphs, and the evaluation of two-point integrals would become considerably more difficult.
@@ 3.2 Three-point correlation functions and the beta functions
The determination of $Z_{h},Z_{g}$-factors is quite similar as that of $Z_{\varphi},Z_{\sigma}$-factors. We start from OPE and arrive at the UV finiteness conditions as eqn.(25), (26). Once the large momentum expansion of the bare correlation functions are computed, we can determine $Z_{g}$ and $Z_{h}$ from the UV finiteness conditions. Remind that the UV divergence is computed by all possible three-point topologies with one leg removed. Following the computation strategy, the one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(3), as well as their independent cut graphs.
@@
The $Z_{g}$-factor is determined by $\langle\varphi\bar{\varphi}\sigma\rangle$ with a $\sigma$ field removed, and the $Z_{h}$-factor is determined by $\langle\sigma\sigma\sigma\rangle$ with a $\sigma$ field removed. The symmetry factors and coupling factors of cut graphs are listed in Table.(1).
@@
The UV divergence is computed by TwoPoint command for two-point integral of cut graphs, with the results [^4] Again an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown,
@@
Special attention should be paid to the tree-level contribution, which is evaluated by the two-point integral of a three-point vertex with one leg removed. Evaluated in the coordinate space, the result is not 1, but
@@
Appropriate $\epsilon$ order terms should be included for the $L$-loop renormalization computation. Then the bare correlation functions for two-loop computation are readily obtained as [^5] For three-point function, $n_{p}=3$. As discussed in the two-point case, each loop order has a factor i-i, while at the tree-level the factor is also i-i. If factoring out an overall factor i-i, the three-point correlation function can be evaluated by tree contribution plus all loop contributions.,
@@
where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factors, and $g_{\Phi}=g$ for $\langle\varphi\bar{\varphi}\sigma\rangle$ while $g_{\Phi}=h$ for $\langle\sigma\sigma\sigma\rangle$.
The UV finiteness conditions of eqn.(25) and eqn.(26) can be realized by,
@@
After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get equations for $Z$-factors by the vanishing of $\epsilon$ poles in each $\ln\frac{\mu^{2}}{p^{2}}$ term [^6] The bare coupling in the denominator will introduce a $\mu^{\frac{\epsilon}{2}}$, which cancels the $\mu^{\frac{\epsilon}{2}}$ factor in numerator. In the bare correlation function, at $L$-loop order there will be a $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor from integration, and a $\mu^{\frac{\epsilon}{2}(2L+1)}$ factor from the couplings. If we take out an overall factor $\mu^{\frac{\epsilon}{2}}$ which will not affect the $\epsilon$ expansion, then each $L$-loop is assigned with a $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor contributing to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion at $\epsilon\to 0$ limit..
For $Z_{g}$-factor, at one-loop there is only one equation from the vanishing of $1/\epsilon$ pole of the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term, which yields,
@@
At two-loop order, there will be two equations from the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term,
@@
and one equation from the sub-Log term,
@@
Again we confirm that the one-loop $Z$-factors can be solved as a whole by the equation of sub-Log term. Consequently, from the equations of leading Log terms we get
@@
Similarly, solving the equations of UV finiteness conditions for $Z_{h}$-factors, we determine the two-loop $Z_{h}$-factor. We get $z_{h}^{11}=-(g^{3}N+h^{3})/h$, and
@@
The higher-loop $Z_{h},Z_{g}$-factors can be similarly determined by including more graphs and more orders of $\epsilon$ series of two-point integrals.
With knowledge of $Z_{h},Z_{g},Z_{\phi}$ and $Z_{\sigma}$, we can compute beta functions by solving equations (9). After expanding the solutions at the limits $g,h\to 0$, we get two-loop results as,
@@
The 2-loop anomalous dimension of field $\varphi$ or $\sigma$ is then given by definition (10) as,

### v4

@@ 3 Two-loop demonstration
For illustrative purposes, this section presents detailed computations of the two-loop anomalous dimensions of the $\varphi^{Q}$ operator. Beta functions and anomalous dimensions are derived from $Z$-factors following standard quantum field theory definitions, where the $Z$-factors themselves are determined via the OPE based algorithm. This algorithm provides a non-trivial implementation of the connection between OPE Wilson coefficients and anomalous dimensions, enabling the systematic determination of $Z$-factors through summation over two-point integrals. The HyperlogProcedures package HP developed from graphical function method assures the computation of two-point integral to very high loop order. The computational work-flow is formalized as a sequential algorithm, proceeding through four key steps: (1) computing the symmetry factors of topologies, (2) evaluating the coupling factors of Feynman diagrams, (3) determining the UV divergences of two-point integrals, and (4) extracting $Z$-factors from UV finiteness conditions.
@@ 3.1 Two-point correlation functions and the field renormalization
We begin with the computation of field renormalization. As specified in eqn.(21), the $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ are derived from the two-point correlation functions $\langle\varphi\bar{\varphi}\rangle$ and $\langle\sigma\sigma\rangle$, respectively. Notably, while the derivations coincide with those of traditional renormalization methods, we frame it explicitly within the OPE-based algorithm for consistency. The one-loop and two-loop topologies in the $\Phi$-field representation are illustrate in Fig.(2).
@@
Each topology represents a set of diagrams with distinct valid internal configurations of the $\sigma$ and $\varphi_{i}$ fields, where these sets are characterized by their respective coupling factors. The symmetry factors and coupling factors for these topologies are listed in the table below,
@@
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$ 1/2 $h^{2}+Ng^{2}$ $2g^{2}$
$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$ 1/2 $N(2g^{4}+g^{2}h^{2})+h^{4}$ $Ng^{4}+2g^{4}+g^{2}h^{2}$
$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ 1/2 $N(g^{4}+2g^{3}h)+h^{4}$ $2g^{4}+2g^{3}h$
No external legs need to be cut from these topologies, and their UV divergences are computed directly using the TwoPoint command in the HyperlogProcedures package. Feynman integrals in the $\overline{\mbox{MS}}$ dimensional regularization scheme are evaluated as power series in $\epsilon$, with results presented below (an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral is implicitly included and not explicitly shown),
@@
The bare correlation functions are then obtained as
@@
where $\Phi$ corresponds to either $\sigma$ or $\varphi$ in the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$. Care must be taken with the sign of each term in the summation. Bare correlation functions include contributions from all tree-level and loop Feynman diagrams, but the TwoPoint command only computes the scalar integrals, omitting $i$-factors from vertices, propagators, or other possible sources. These $i$-factors must be recovered using Feynman rules derived for the Euclidean-space Lagrangian. Graphical analysis shows the overall factor is $(-1)^{n_{p}}i$, where $n_{p}$ is the number of external legs. For two-point integrals ($n_{p}=2$), each loop order carries a factor $i$, while the tree-level contrition has a factor i-i. Factoring out an overall i-i, the two-point correlation function is expressed as the tree-level contribution minus the sum of loop contributions.
@@
The renormalized two-point correlation functions eqn.(21) must be UV finite, leading to the UV finiteness conditions
$Z_{\varphi}G(g_{0},h_{0};\varphi(x_{0}),\bar{\varphi}(x_{1}))=\mbox{UV~finite}~~~,~~~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{0}),\sigma(x_{1}))=\mbox{UV~finite}~.$ (42)
@@
To derive these conditions explicitly, firstly we rewrite the bare couplings $g_{0},h_{0}$ in the bare correlation functions in terms of renormalized couplings $g,h$ using eqn.(7), then expand the full expression in the double limit $g,h\to 0$, followed by $\epsilon\to 0$. Note that eqn.(7) implies each $g$ or $h$ is accompanied by a factor $\mu^{\frac{\epsilon}{2}}$, so $L$-loop contributions carry a $\mu^{L\epsilon}$ factor. Combined with the implicit $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor, $L$-loop terms include an overall $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor. Expanding this factor around $\epsilon\to 0$ generates terms proportional to powers of $\ln\frac{\mu^{2}}{p^{2}}$. Since each power of $\ln\frac{\mu^{2}}{p^{2}}$ is independent, UV finiteness requires the coefficient of every $\ln\frac{\mu^{2}}{p^{2}}$ term to be free of $\epsilon$ poles, introducing additional equations beyond those from the leading terms. Inspecting the UV divergence structures of $L$-loop integrals, the leading divergence is $1/\epsilon^{L}$}, so the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ order generally yields $L$ equations for canceling $L$ poles. For $\ln^{k}\frac{\mu^{2}}{p^{2}}$ terms, the leading divergence is $1/\epsilon^{L-k}$, producing (Lk)(L-k) equations. We may either use only equations from leading Log terms [^3] Leading Log terms refer to those with $\ln^{0}(\mu^{2}/p^{2})$ dependence, sub-Log terms to $\ln(\mu^{2}/p^{2})$ dependence, and so on., or include those from higher Log orders. If only leading Log equations are used, $Z_{\varphi},Z_{\sigma},Z_{g},Z_{h}$ must be solved simultaneously to fully determine the $Z$-factors. In contrast, incorporating sub-Log equations allows each $Z$-factor to be determined independently via it own UV finiteness conditions.
@@
We now analyze the UV finiteness conditions for the $Z_{\varphi}$-factor. At one-loop order, following the expansions outlined above, we obtain
@@
The coefficients of the sub-Log terms are already UV finite and do not yield additional equations. For the leading Log term to be finite, the $1/\epsilon$ pole must cancel, leading to the condition,
@@
At two-loop order, we solve for the coefficients $z^{21}_{\varphi},z^{22}_{\varphi}$. Expanding the expression, the cancelation of $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term gives two equations,
@@
while the cancelation of the $1/\epsilon$ pole in the sub-Log term provides a third equation,
@@
Computations can be performed iteratively by loop order, solving for $Z$-factors using only leading Log term equations. For instance, using the two leading Log equations above, determining $z_{\varphi}^{21}$ and $z_{\varphi}^{22}$ requires prior knowledge of one-loop $Z$-factors, including $Z_{g},Z_{h},Z_{\varphi},Z_{\sigma}$. Alternatively, incorporating UV finiteness conditions from all Log terms of a correlation function generates a sufficient set of equations to solve for the corresponding $Z$-factor independently, without relying on lower-loop $Z$-factors from other channels. For instance, the sub-Log term equation can be solved to obtain a combination of one-loop $Z$-factors. This combination directly appears in the leading Log equations. Substituting this result back into the two leading Log equations fully determines the two-loop $Z_{\varphi}$-factor. As a byproduct, sub-Log term equations also provide consistency checks for lower-loop $Z$-factors computed from other correlation functions.
Following this approach, we determine the two-loop $Z_{\varphi}$-factor coefficients as
@@
The $Z_{\sigma}$-factor is determined via an identical procedure, with the only difference being the coupling factors. After analogous computations, the two-loop $Z_{\sigma}$-factor coefficients are
@@
Higher-loop $Z_{\varphi}$ and $Z_{\sigma}$-factor can be determined using the same iterative procedure. However, this becomes increasingly computationally demanding due to the exponential growth in the number of Feynman diagrams and the heightened complexity of evaluating two-point integrals at higher loops.
@@ 3.2 Three-point correlation functions and the beta functions
We now apply the proposed algorithm to determine the $Z_{h}$ and $Z_{g}$-factors. Starting from the corresponding OPEs, we derive the UV finiteness conditions given in eqn.(25) and (26). These equations non-trivially relate the respective $Z$-factor ratios to the sum of two-point integrals, which are generated by removing one external leg from three-point topologies in all possible ways. Following the algorithmic strategy, the one-loop and two-loop topologies in the $\Phi$-field representation, along with their independent cut graphs, are illustrated in Fig.(3). The $Z_{g}$-factor is determined from $\langle\varphi\bar{\varphi}\sigma\rangle$ with one $\sigma$ field removed, while the $Z_{h}$-factor is determined from $\langle\sigma\sigma\sigma\rangle$ with one $\sigma$ field removed. The symmetry factors and coupling factors of the cut graphs are listed in Table.(1).  [^4] We note that the symmetry factor defined herein is given by (n1)!(n-1)! divided the internal symmetry of the diagram and the permutation symmetry of external legs, where $n$ is the number of external legs. This differs from the conventional definition, which only accounts for internal symmetry, as used in Fei:2014xta . For example, using data from Table.(1), the conventional symmetry factors for two-loop topologies are $S(\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}}=\frac{1}{2}$, $S(\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}}=1$, $S(\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=\frac{1}{2}$. The ratios of our symmetry factors to the conventional ones are $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})/S(\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=2$ and $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})/S(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=2$. These differences can be absorbed into the coupling factors if the conventional symmetry factor definition is adopted, effectively encoding the summation of external legs into structure factors.
UV divergences for the two-point integrals of the cut graphs are computed using TwoPoint command in HyperlogProcedures package. The one-loop results [^5] As before, an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral is implicit included and not explicitly shown. are
@@
Special care is required for the tree-level contribution, which is evaluated from the two-point integral of a three-point vertex with one leg removed. In coordinate space, this result is not unity but,
@@
For $L$-loop renormalization computations, the appropriate orders of $\epsilon$ must be included. The bare correlation functions for two-loop computations are then obtained as [^6] For three-point function, $n_{p}=3$, so the overall $i$-factors for tree and loop contributions have the same sign.
@@
where $\Phi$ corresponds to $\sigma$ or $\varphi$ in the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$, and $g_{\Phi}=g$ for $\langle\varphi\bar{\varphi}\sigma\rangle$ while $g_{\Phi}=h$ for $\langle\sigma\sigma\sigma\rangle$.
The UV finiteness conditions of eqn.(25) and (26) are explicitly,
@@
After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ using relations eqn.(7), and expanding the full expression in the double limit $g,h\to 0$, followed by $\epsilon\to 0$, we obtain equations for $Z$-factors by enforcing the cancelation of $\epsilon$ poles in each $\ln\frac{\mu^{2}}{p^{2}}$ term [^7] The bare coupling in the denominator introduces a $\mu^{\frac{\epsilon}{2}}$, which cancels the $\mu^{\frac{\epsilon}{2}}$ factor in the numerator. In the bare correlation function, $L$-loop contribution include a $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor from integration and a $\mu^{\frac{\epsilon}{2}(2L+1)}$ factor from couplings. Factoring our an overall $\mu^{\frac{\epsilon}{2}}$ (which does not affect the $\epsilon$ expansion), each $L$-loop term carries a $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor that contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion as $\epsilon\to 0$..
For the $Z_{g}$-factor, the one-loop equation yields,
@@
At two-loop order, the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term provides two equations,
@@
while the sub-Log term provides a third equation,
@@
As before, the sub-Log term equation allows solving for the combination of one-loop $Z$-factors. Substituting this result into the leading Log equations fully determines the two-loop $Z_{g}$-factor coefficients,
@@
Following analogous steps for the $Z_{h}$-factors, we solve its UV finiteness conditions to obtain the two-loop results. The one-loop coefficient is $z_{h}^{11}=-(g^{3}N+h^{3})/h$, and the two-loop coefficients are
@@
Higher-loop $Z_{h}$ and $Z_{g}$-factors can be determined similarly by including additional topologies and higher-order $\epsilon$ terms in the two-point integrals.
With $Z_{h},Z_{g},Z_{\phi}$ and $Z_{\sigma}$ determined, the beta functions are computed by solving eqn.(9). Expanding the solutions in the limit $g,h\to 0$, the two-loop results are
@@
The two-loop anomalous dimension of the $\varphi$ and $\sigma$ fields are then obtained via eqn.(10),